OpenStudy (friendzone):

What is the slope of the line segment graphed here? A) -1 Eliminate B) 0 C) 1 D) 5

OpenStudy (anonymous):

where is the graph

OpenStudy (anonymous):

anyway the slope= (y2-y1)/(x2-x1)

OpenStudy (anonymous):

so the slope =(2-2)/(6-1)=0/5=0